OpenStudy (anonymous):

An isoceles triangle is drawn with base angels 32 and base 24cm. Find the base angles of an isosceles triangle with the same base but will double the area.

OpenStudy (nincompoop):

first, learn what isosceles triangle is

OpenStudy (anonymous):

have you learned the cosine formula? c^2=a^2+b^2 -2(a)(b)cos(x) ?

OpenStudy (anonymous):

oh way no need we can simply use tan(x)=opp/adj which is tan(32)=leg/12 and you find the leg are you familiar with this process?

OpenStudy (anonymous):

wait*

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

i got it until there.... what about the rest

OpenStudy (anonymous):

give me a sec ill picture the detailed answer

OpenStudy (anonymous):

see

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

yes i got it thanks

OpenStudy (anonymous):

Refer to the plot and solution from the Mathematica 9 program.